Required information

[The following information applies to the questions displayed below.]

Mason Corporation began operations at the beginning of the current year. One of the company’s products, a refrigeration element, sells for $185 per unit. Information related to the current year’s activities follows.

Variable costs per unit:
Direct material $ 15
Direct labor 35
Manufacturing overhead 44
Annual fixed costs:
Manufacturing overhead $ 600,000
Selling and administrative 860,000
Production and sales activity:
Production (units) 24,000
Sales (units) 20,000

Mason carries its finished goods inventory at the average unit cost of production and is subject to a 30 percent income tax rate. There was no work in process at year-end.

1. Determine the cost of the December 31 finished goods inventory.

2. Compute Mason’s net income for the current year ended December 31

3. If next year’s production decreases to 23,000 units and general cost behavior patterns do not change, what is the likely effect on:

a. The direct-labor cost of $35 per unit?

  • No change
  • Increase
  • Decrease

b. The fixed manufacturing overhead cost of $600,000?

  • No change
  • Increase
  • Decrease

c.  The fixed selling and administrative cost of $860,000?

  • No change
  • Increase
  • Decrease

d. The average unit cost of production?

  • No change
  • Increase
  • Decrease

.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Mason Corporation
Calculation of Average costs per unit Amount $ Note
Variable costs
Direct material                 15.00
Direct labor                 35.00
Manufacturing overhead                 44.00
Variable costs per unit                 94.00 A
Production (units)         24,000.00 B
Variable costs of Production 2,256,000.00 C=A*B
Add: Annual fixed costs:
Manufacturing overhead       600,000.00 D
Total costs of Production 2,856,000.00 E=C+D
Production (units)         24,000.00 See B
Average costs per unit               119.00 F=E/B
Answer 1
Production (units)         24,000.00 See B
Sales (units)         20,000.00 G
Ending inventory           4,000.00 H=B-G
Average costs per unit               119.00 See F
Cost of Dec 31 Finished goods inventory       476,000.00 I=H*F
Sales (units)         20,000.00 See G
Average costs per unit               119.00 See F
Cost of goods sold 2,380,000.00 J=G*F
Sales (units)         20,000.00 See G
Sell price per unit               185.00 K
Total sales 3,700,000.00 L=G*K
Answer 2
Income Statement Amount $
Sales    3,700,000.00 See L
Less: Cost of goods sold    2,380,000.00 See J
Gross Profit 1,320,000.00
Less: Selling and administrative       860,000.00
Income before tax       460,000.00 M
Less: Income tax       138,000.00 N=M*30%
Net Income       322,000.00 O=M-N
Answer 3 a
Variable costs per unit does not change with change in activity unless otherwise mentioned. Here it is given that next year also general cost behavior patterns will not change, so:
The direct-labor cost of $35 per unit will be same and there will be no change.
Answer 3 b
Total fixed costs does not change with change in activity unless otherwise mentioned. Here it is given that next year also general cost behavior patterns will not change, so:
The fixed manufacturing overhead cost of $600,000 will be same and there will be no change.
Answer 3 c
Total fixed costs does not change with change in activity unless otherwise mentioned. Here it is given that next year also general cost behavior patterns will not change, so:
The fixed selling and administrative cost of $860,000 will be same and there will be no change.
Answer 3 d
Calculation of average unit cost of production Amount $
Variable costs
Direct material                 15.00
Direct labor                 35.00
Manufacturing overhead                 44.00
Variable costs per unit                 94.00 See A
Production (units)         23,000.00 P
Variable costs of Production 2,162,000.00 Q=A*P
Add: Annual fixed costs:
Manufacturing overhead       600,000.00 See D
Total costs of Production 2,762,000.00 R
Production (units)         23,000.00 See P
Average costs per unit               120.09 S=R/P
Increase by $               1.09 T=S-F
So average unit cost of production will increase by $ 1.09.